# Stochastic Processes and Transport Plans

### v2

@@ 3. Finally, we conclude
We recap the example in [27, Example II] with our approach.
@@
By Corollary 5.10, we obtain that $(\mathcal{K}_{t}(y,\cdot),v^{y}_{t}$ with the velocity field
@@ 6.1 Conditional Expectation
Conditional expectations $\mathbb{E}[X|Y]$ of random vectors can also be interpreted as Borel measurable functions $\psi:\mathbb{R}^{d}\to\mathbb{R}^{d}$ via the Doob–Dynkin Lemma, see [22, Lemma 1.14].
@@ 6.2 Curve via Conditional Expectation
In [29, Theorem 3.3] it was proven that for well-behaved stochastic processes $(X_{t})_{t}$ the conditional expectation can be used to obtain a vector field $v_{t}$ such that $(X_{t,\sharp}\mathbb{P},v_{t})$ fulfills the continuity equation. For convenience, we include the proof.
@@
Let $X_{t}=(1-t)X_{0}+tX_{1}$ for independent random variables $X_{0},X_{1}$, where $X_{0}\sim\mathcal{N}(0,I_{d})$. Then in [43, Appendix A1] it is shown that the formula for the vector field associated with $\mu_{t}$,
@@
from Proposition 4.11 can also be derived via the velocity field from Theorem 6.3 and Tweedie’s formula. Recall that Tweedie’s formula [12, Lemma 3.2] for $X_{t}$ reads as
@@ 7 Flow Matching
Let $\mu_{1}\in\mathcal{P}_{2}(\mathbb{R}^{d})$, and let $\mathcal{K}_{t}(y,\cdot):[0,1]\to\mathcal{P}_{2}(\mathbb{R}^{d})$ be absolutely continuous curves with associated velocity fields $v_{t}^{y}$ for $\mu_{1}$-a.e. $y\in\mathbb{R}^{d}$. Further, assume that $(t,x,y)\mapsto v_{t}^{y}$ is measurable and $\|v_{t}^{y}\|_{L^{2}(\mathbb{R}^{d}\times\mathbb{R}^{d},\alpha_{t})}\in L^{1}([0,1])$. For $\alpha_{t}\coloneqq\mathcal{K}_{t}(y,\cdot)\times_{y}\mu_{1}$, we consider the absolutely continuous curve $\nu_{t}\coloneqq\pi^{1}_{\sharp}\alpha_{t}.$ Then its associated velocity field $v:[0,1]\times\mathbb{R}^{d}\to\mathbb{R}^{d}$ defined by (74) fulfills
@@
$\mathbb{E}_{(t,x,y)\sim\alpha_{t}\times_{t}\mathcal{L}_{(0,1)}}\left[\|u_{t}(x)\|^{2}\right]$ $=\mathbb{E}_{t\sim\mathcal{L}_{(0,1)}}\mathbb{E}_{y\sim\nu_{t}}\left[\mathbb{E}_{y\sim\alpha_{t}^{x}}\left[\|u_{t}(x)\|^{2}\right]\right]$ (146)
@@
$=\mathbb{E}_{t\sim[0,1]}\mathbb{E}_{y\sim\mu_{1}}\mathbb{E}_{x\sim\mathcal{N}(0,I_{d})}\big{[}\|u_{t}((1-tr)x+ty)-(y-rx)\|^{2}\big{]}$ (156)
$=\mathbb{E}_{t\sim[0,1],(x,y)\sim\mu_{0}\times\mu_{1}}\big{[}\|u_{t}((1-tr)x+ty)-(y-rx)\|^{2}\big{]}.$ (157)
@@
$\mathbb{E}[X|Y={\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\cdot}]=\operatorname*{argmin}_{f\in L^{2}(\mathbb{R}^{d},Y_{\sharp}\mathbb{P})}\mathbb{E}_{\mathbb{P}}[\|X-f(Y)\|^{2}].$
@@
$v$ $=\operatorname*{argmin}_{u\in L^{2}(\mathbb{R}^{d},\mu_{t}\times\mathcal{L}_{(0,1)})}\mathbb{E}_{\mathcal{L}_{(0,1)}\times\mathbb{P}}[\|\partial_{t}X_{t}-u_{t}(X_{t})\|^{2}].$ (163)
@@
$v$ $=\operatorname*{argmin}_{u\in L^{2}(\mathbb{R}^{d},\mu_{t}\times\mathcal{L}_{(0,1)})}\mathbb{E}_{\mathcal{L}_{(0,1)}\times\mathbb{P}}[\|X_{1}-X_{0}-u_{t}(X_{t})\|^{2}]$ (164)
$=\operatorname*{argmin}_{u\in L^{2}(\mathbb{R}^{d},\mu_{t}\times\mathcal{L}_{(0,1)})}\mathbb{E}_{(t,x,y)\sim\mathcal{L}_{(0,1)}\times P_{X_{0},X_{1}}}[\|y-x-u_{t}(e_{t}(x,y))\|^{2}]$ (165)
@@ 8.2 Numerical Examples
Finally, we show a high dimensional application of flow matching in Figure 8. Here we used the flow vector field from [30], which was already trained via minibatch OT flow matching on images of cat faces. For a comparison of the performance for the flow matching with different plans see [40, Table 2].
@@ 9.1 Conditional Wasserstein Distance
The notion of conditional Wasserstein distances appeared already in [25, Section 2.4] as well as in the PhD thesis of Gigli [14, Chapter 4], who studied a metric on the so-called ”geometric tangent space” of the Wasserstein-$2$ space. Recently, conditional Wasserstein distances were studied in the realm of machine learning in [4] as well as in functional optimal transport in [21]. We will focus on its application in Bayesian flow matching as presented in our paper [7], where also the proofs of the following theorems and propositions are provided.
@@
As shown in [23, Theorem 3,4] and [4], there is a counterpart of Theorem 3.3 for the conditional Wasserstein distance.
@@ 9.3 Bayesian Flow Matching
ii) We can use Proposition 9.1 ii) and solve the optimal transport problem for fixed samples $w$ from $\eta$, to obtain an optimal coupling $\alpha\in\Gamma_{\eta,o}(\mu_{0},\mu_{1})$. In order to ensure that the $\alpha_{w}\in\Gamma_{o}(\mu_{0}^{w},\mu_{1}^{w})$ can be used to build $\alpha\in\Gamma_{\eta,o}(\mu_{0},\mu_{1})$, some measurability conditions must be ensured. For examples where these conditions are fulfilled, see e.g. [15, Corollary 1.2] and [7, Proposition 8].
@@ 9.4 Numerical Examples of Bayesian Flow Matching
The noise variable $\Xi$, which is independent from $X$, is distributed as $\mathcal{N}(0,0.1)$. Then by [18, Lemma 11] the posterior distribution $P_{X|Y=y}$ is a Gaussian mixture model which can be analytically computed. This allows us to sample from $P_{X|Y=y}$.
@@ 10.1 Curves of Probability Measures from ODEs
We start with classical results from the theory of ODEs, for existence and regularity see, e.g., [39, Corollary 2.6, Theorem 2.10].
@@ 11 A Glimpse at Score-Based Diffusion
Usually, we do not have access to $\nabla\log p_{t}$. Fortunately, by [41, Appendix], the loss function can be rewritten, up to a constant, as

### v4

@@ 3. Finally, we conclude
We recap the example in [28, Example II] with our approach.
@@
By Corollary 5.10, we obtain that $(\mathcal{K}_{t}(y,\cdot),v^{y}_{t})$ with the velocity field
@@ 6.1 Conditional Expectation
Conditional expectations $\mathbb{E}[X|Y]$ of random vectors can also be interpreted as Borel measurable functions $\psi:\mathbb{R}^{d}\to\mathbb{R}^{d}$ via the Doob–Dynkin Lemma, see [23, Lemma 1.14].
@@ 6.2 Curve via Conditional Expectation
In [30, Theorem 3.3] it was proven that for well-behaved stochastic processes $(X_{t})_{t}$ the conditional expectation can be used to obtain a vector field $v_{t}$ such that $(X_{t,\sharp}\mathbb{P},v_{t})$ fulfills the continuity equation. For convenience, we include the proof.
@@
Let $X_{t}=(1-t)X_{0}+tX_{1}$ for independent random variables $X_{0},X_{1}$, where $X_{0}\sim\mathcal{N}(0,I_{d})$. Then in [45, Appendix A1] it is shown that the formula for the vector field associated with $\mu_{t}$,
@@
from Proposition 4.11 can also be derived via the velocity field from Theorem 6.3 and Tweedie’s formula. Recall that Tweedie’s formula [11, Lemma 3.2] for $X_{t}$ reads as
@@ 7 Flow Matching
Let $\mu_{1}\in\mathcal{P}_{2}(\mathbb{R}^{d})$, and let $\mathcal{K}_{t}(y,\cdot):[0,1]\to\mathcal{P}_{2}(\mathbb{R}^{d})$ be absolutely continuous curves with associated velocity fields $v_{t}^{y}$ for $\mu_{1}$-a.e. $y\in\mathbb{R}^{d}$. Further, assume that $(t,x,y)\mapsto v_{t}^{y}$ is measurable and $\|v_{t}^{y}\|_{L^{2}(\mathbb{R}^{d}\times\mathbb{R}^{d},\alpha_{t})}\in L^{2}([0,1])$. For $\alpha_{t}\coloneqq\mathcal{K}_{t}(y,\cdot)\times_{y}\mu_{1}$, we consider the absolutely continuous curve $\nu_{t}\coloneqq\pi^{1}_{\sharp}\alpha_{t}.$ Then its associated velocity field $v:[0,1]\times\mathbb{R}^{d}\to\mathbb{R}^{d}$ defined by (81) fulfills
@@
$\mathbb{E}_{(t,x,y)\sim\alpha_{t}\times_{t}\mathcal{L}_{(0,1)}}\left[\|u_{t}(x)\|^{2}\right]$ $=\mathbb{E}_{t\sim\mathcal{L}_{(0,1)}}\mathbb{E}_{x\sim\nu_{t}}\left[\mathbb{E}_{y\sim\alpha_{t}^{x}}\left[\|u_{t}(x)\|^{2}\right]\right]$ (153)
@@
$=\mathbb{E}_{t\sim[0,1]}\mathbb{E}_{y\sim\mu_{1}}\mathbb{E}_{x\sim\mathcal{N}(0,I_{d})}\big{[}\|u_{t}^{\theta}((1-tr)x+ty)-(y-rx)\|^{2}\big{]}$ (163)
$=\mathbb{E}_{t\sim[0,1],(x,y)\sim\mu_{0}\times\mu_{1}}\big{[}\|u_{t}^{\theta}((1-tr)x+ty)-(y-rx)\|^{2}\big{]}.$ (164)
@@
$\mathbb{E}[X|Y=\cdot]=\operatorname*{argmin}_{f\in L^{2}(\mathbb{R}^{d},Y_{\sharp}\mathbb{P})}\mathbb{E}_{\mathbb{P}}[\|X-f(Y)\|^{2}].$
@@
$v$ $=\operatorname*{argmin}_{u\in L^{2}(\mathbb{R}^{d},\mu_{t}\times_{t}\mathcal{L}_{(0,1)})}\mathbb{E}_{\mathcal{L}_{(0,1)}\times\mathbb{P}}[\|\partial_{t}X_{t}-u_{t}(X_{t})\|^{2}].$ (170)
@@
$v$ $=\operatorname*{argmin}_{u\in L^{2}(\mathbb{R}^{d},\mu_{t}\times_{t}\mathcal{L}_{(0,1)})}\mathbb{E}_{\mathcal{L}_{(0,1)}\times\mathbb{P}}[\|X_{1}-X_{0}-u_{t}(X_{t})\|^{2}]$ (171)
$=\operatorname*{argmin}_{u\in L^{2}(\mathbb{R}^{d},\mu_{t}\times_{t}\mathcal{L}_{(0,1)})}\mathbb{E}_{(t,x,y)\sim\mathcal{L}_{(0,1)}\times P_{X_{0},X_{1}}}[\|y-x-u_{t}(e_{t}(x,y))\|^{2}]$ (172)
@@ 8.2 Numerical Examples
Finally, we show a high dimensional application of flow matching in Figure 8. Here we used the flow vector field from [31], which was already trained via minibatch OT flow matching on images of cat faces. For a comparison of the performance for the flow matching with different plans see [42, Table 2].
@@ 9.1 Conditional Wasserstein Distance
Conditional Wasserstein distances appear already in [26, Section 2.4], [32] as well as in the PhD thesis of Gigli [13, Chapter 4], who studied a metric on the so-called ”geometric tangent space” of the Wasserstein-$2$ space. Recently, conditional Wasserstein distances were studied in the realm of machine learning in [4] as well as in functional optimal transport in [21]. We will focus on its application in Bayesian flow matching as presented in our paper [7], where also the proofs of the following theorems and propositions are provided.
@@
As shown in [24, Theorem 3,4] and [4], there is a counterpart of Theorem 3.3 for the conditional Wasserstein distance.
@@ 9.3 Bayesian Flow Matching
ii) We can use Proposition 9.1 ii) and solve the optimal transport problem for fixed samples $w$ from $\eta$, to obtain an optimal coupling $\alpha\in\Gamma_{\eta,o}(\mu_{0},\mu_{1})$. In order to ensure that the $\alpha_{w}\in\Gamma_{o}(\mu_{0}^{w},\mu_{1}^{w})$ can be used to build $\alpha\in\Gamma_{\eta,o}(\mu_{0},\mu_{1})$, some measurability conditions must be ensured. For examples where these conditions are fulfilled, see e.g. [14, Corollary 1.2] and [7, Proposition 8].
@@ 9.4 Numerical Examples of Bayesian Flow Matching
The noise variable $\Xi$, which is independent from $X$, is distributed as $\mathcal{N}(0,0.1)$. Then by [17, Lemma 11] the posterior distribution $P_{X|Y=y}$ is a Gaussian mixture model which can be analytically computed. This allows us to sample from $P_{X|Y=y}$.
@@ 10.1 Curves of Probability Measures from ODEs
We start with classical results from the theory of ODEs, for existence and regularity see, e.g., [41, Corollary 2.6, Theorem 2.10].
@@ 11 A Glimpse at Score-Based Diffusion
Usually, we do not have access to $\nabla\log p_{t}$. Fortunately, by [43, Appendix], the loss function can be rewritten, up to a constant, as
@@
Acknowledgement. Many thanks to R. Beinert and R. Duong for reading the manuscript and to R. Duong for pointing out Example 4.12. C. Wald and G. Steidl gratefully acknowledge funding by the DFG within the SFB “Tomography Across the Scales” (STE 571/19-1, project number: 495365311) and G. Steidl acknowledges funding by the Deutsche Forschungsgemeinschaft under Germany’s Excellence Strategy (The Berlin Mathematics Research Center MATH+).